Suppose a teacher recorded the attendance of her students in a recent statistics class because she wanted to investigate the Missed Classes Final Grade
4 75
6 80
1 91
4 71
0 95
2 83
0 91
5 86
2 96
Answer #1

Hese, data is, the given missed Final classes(y) arrades (9) 4 75 X4 ŷ 300 2 (y-ý) 44.2185 15-0288 2 6 80 480 81.6497 76.1233bo ý -5, - 85333 - (-24632) ( 2.66+) = 85.333 + 7.3695 5o = 92.7025 so - 50+5 = 927025 2.7632x Here, sse - 2 cy-ŷ)? ose 327.83 The statistic is, Lest E bı - Bi sech) h Sech) 2.7682 1.1102 2-48892 Ans. t = -2.49 4 P. value p.value -p [I Łl >-2-49) = 02.7632 - 3.885) 2.7632 t 8.8851) (-2.1 - 2-7632 +3.885)) 커 (-6 -6.648, 1.122 Ans. confidence Interval is, - 6.6483, 1. 1219 -